Question 2. (20 pts.) a) Simplify 2 (2< (51/2)^(2 < (-1/3) 2<(-1/6) b) Solve 24 + Z2 + 1 = 0
Answer #1

a (2<(**))*(2<4- =))} 2) = De. Moirre theory 4 <2x s):24 245) = 2 <E) il < (2*4= 215 +1) =8< ( SA 22 5 + 2 =8<120) = (Auswa)